OpenStudy (anonymous):

P=2L+2W for W i came up with W=(P-2L)2 is this correct?

OpenStudy (anonymous):

P=2L+2w P-2L=2L-2L+2w P-2L=0+2w (p-2L)/2=2W/2 (P-2L)/2=W W=(P-2L)/2 final answer

OpenStudy (anonymous):

Looks good.

OpenStudy (anonymous):

Thanks1

OpenStudy (anonymous):

Sure thing :)